# Comprehensive Study on German Language Models for Clinical and Biomedical Text Understanding

### v1

@@ Abstract
| Ahmad Idrissi-Yaghir, Amin Dada, Henning Schäfer, |
| Kamyar Arzideh, Giulia Baldini, Jan Trienes, Max Hasin, |
| Jeanette Bewersdorff, Cynthia S. Schmidt, Marie Bauer, |
| Kaleb E. Smith, Jiang Bian, Yonghui Wu, Jörg Schlötterer, |
| Torsten Zesch, Peter A. Horn, Christin Seifert, |
| Felix Nensa, Jens Kleesiek,Christoph M. Friedrich |
| Department of Computer Science, University of Applied Sciences and Arts Dortmund, Dortmund, Germany |
| Institute for Medical Informatics, Biometry and Epidemiology, University Hospital Essen, Essen, Germany |
| Institute for AI in Medicine (IKIM), University Hospital Essen (AöR), Essen, Germany |
| Institute for Transfusion Medicine, University Medicine Essen, Essen, Germany |
| Computational Linguistics, CATALPA FernUniversität in Hagen, Germany |
| NVIDIA, Santa Clara, CA, USA |
| Department of Health Outcomes and Biomedical Informatics, College of Medicine, |
@@
| Cancer Informatics and eHealth core, University of Florida Health Cancer Center, |
@@
| University of Marburg, Marburg, Germany |
| University of Mannheim, Mannheim, Germany |
| University Hospital Essen, Institute of Interventional and Diagnostic Radiology and Neuroradiology, |
@@
| Cancer Research Center Cologne Essen (CCCE), West German Cancer Center Essen, |
@@
| German Cancer Consortium (DKTK, Partner site Essen), Heidelberg, Germany |
| Department of Physics, TU Dortmund, Dortmund, Germany |
@@ 1. Introduction
000 These authors contributed equally to this workIn recent years, pre-trained language models (PLMs) such as BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019) have become crucial in the field of natural language processing (NLP). These models have significantly enhanced the performance of a wide range of tasks, including document and token classification, as well as machine translation and text summarization. The success of these models is primarily attributed to their transformer-based architecture Vaswani et al. (2017) and their training on large amounts of unlabeled data. However, since these models are often trained on general data sources such as Wikipedia, news articles, and books, their effectiveness may be limited in specific domains, such as medicine or finance, which have distinct terminologies and writing styles. To achieve better results in these specialized fields, it is essential to use language models that are tailored to these specific domains. Building on this concept, language models can be adapted to specialized domains through two methods. The first approach involves training models from scratch on unlabeled data from the desired domain. This method ensures that the model is grounded in the unique characteristics of the target domain from the beginning. The second approach relies on continuous pre-training. Instead of starting from scratch, existing general-purpose pre-trained models can be used and refined through further pre-training on the domain-specific unlabeled data Gururangan et al. (2020). This allows a transition that shifts the focal point of the model from a broad scope to one specific to the particularities of the target domain.
@@ 4. Base Models
In this section, the models included in our benchmark are described. As a baseline for our evaluations, we use models that have been pre-trained on extensive datasets and have demonstrated strong performance on a variety of general and medical NLP tasks. The general German language models GBERT, GBERT and GELECTRA Chan et al. (2020) were primarily considered. They were trained on four different datasets: the German portion of the Open Super-large Crawled ALMAnaCH coRpus (OSCAR) (Ortiz Suárez et al., 2020), German Wikipedia dump, The Open Parallel Corpus (OPUS) (Tiedemann, 2012), and Open Legal Data (Ostendorff et al., 2020). Another model is GottBERT (Scheible et al., 2020), a German RoBERTa model trained on the German part of OSCAR data. Furthermore, two multilingual models XLM-R Conneau et al. (2020) and mDeBERTa V3 He et al. (2023) were also considered. Additionally, GeBERTa and GeBERTa Dada et al. (2023) were explored, with these DeBERTa v2 He et al. (2021) based models pre-trained from scratch on a combination of data sources ranging from Wikipedia to medical datasets. Finally, the German biomedical and clinical models BioGottBERT Lentzen et al. (2022) and medBERTde Bressem et al. (2024) were also examined to provide additional comparative insights.
@@ 5. Pre-training
Continuous pre-training of several publicly available general domain models was performed with the aforementioned datasets. First, we continued the pre-training of the GeBERTa and GeBERTa models with the clinical dataset. The objective was to quantify the contribution of clinical data to the performance of the models. While the original model has already been trained on medical texts obtained mainly from MIMIC-III, translated PubMed abstracts, and filtered CC100 Wenzek et al. (2020), it has not seen any real-world German clinical data. Both models were trained for 200k AdamW Loshchilov and Hutter (2019) optimization steps with a batch size of 512. The learning rate was set to 3e for the base model and 2e for the large model. Additionally, the GBERT and GBERT models were further pre-trained on the clinical data with the same parameters. These models were not explicitly pre-trained on medical texts before.
In order to separately quantify the influence of translation-based medical texts and clinical documents, additional pre-training experiments were conducted with the GBERT and GBERT models. Both models underwent further pre-training on the translated dataset. The GBERT was trained for 73k steps with a learning rate of 5e and a batch size of 144. On the other hand, the GBERT was trained for 75k steps using a batch size of 336. Due to hardware limitations, the initial experiments were conducted separately, reflecting the differences in batch sizes between the experiments. The AdamW optimizer was also employed during this pre-training for the optimization process.
@@ 6. Downstream Datasets & Tasks
The RadQA dataset is an extractive question-answering dataset created from 1,223 anonymized radiology reports of brain CT scans from a large hospital in Germany. For its development, three medical student assistants in their sixth and eighth semesters were assigned to annotate 29,273 question-answer pairs. The annotators were provided with a list of questions designed with the input of a radiologist. This decision was influenced by the unique nature of radiology queries and the challenges of annotating sensitive clinical data. The annotators generated one custom question for every third report to ensure variety.
@@ 8. Results
Across all tasks, our clinically pre-trained models achieved the highest F1-Scores, with the exception of GGPONC 2.0, where one of the translation models achieved the highest score. This is especially evident in the results of the GraSCCo dataset, where clinical pre-training improved the performance of GBERT by 5.4 percentage points in F1-Score and the performance of GBERT by 6.9 percentage points. In addition, the GBERT-BioM-translation models were able to outperform the general models. For instance, pre-training on translated texts resulted in an improvement of 2.1 percentage points over GBERT.
Although BRONCO150 is the only NER task with real-world clinical documents, there is no observable performance difference between MedBERTde, GeBERTa, and our clinical base models. However, the GeBERTa-Clinical model outperformed the second-best model by 0.5 percentage points. The minor differences between the various models and the lack of clear advantages of clinical models can be attributed to the small size of the dataset. Interestingly, on GGPONC 2.0, various general domain models are on par with clinical models or even better. In the case of GeBERTa, the additional clinical pre-training decreased its performance. It is worth noting that the dataset consists of guidelines for oncologists written in a different writing style than clinical documents. Overall, the translation-based models achieved the highest results.
In the multi-label classification CLEF eHealth 2019 task, large models generally performed better. For example, GBERT has an F1-Score that is around 1.6 percentage points better than its base variant. Domain-specific continuous pre-training on the translations reached strong performance across all metrics. Prioritizing a balance between precision and recall favors the GELECTRA model, while GeBERTa-Clinical surpasses in precision. The general domain GBERT model provides the highest recall, although its precision is lower compared to others.
Looking at the difference between GBERT-Clinical and GBERT-BioM-Translation in Table 4, the latter achieves comparable results to its clinical counterpart on BRONCO, GGPONC 2.0, and RadQA. Only on GraSCCo the results of the GBERT models trained on the translations are worse. Compared to the baseline GBERT general models, all further pre-trained models have an advantage. In this case, we see no clear advantage of training on German clinical data over translated texts.
@@
| GBERT Chan et al. (2020) | .816 | .818 | .815 | .794 | .707 | .642 | .617 | .676 | |
| GBERT Chan et al. (2020) | .832 | .802 | .865 | .809 | .718 | .647 | .617 | .680 | |
@@
| XLM-R Conneau et al. (2020) | .804 | .781 | .829 | .813 | .731 | .674 | .655 | .694 | |
| GELECTRA Chan et al. (2020) | .827 | .826 | .828 | .812 | .725 | .681 | .702 | .661 | |
| general | mDeBERTa V3 He et al. (2023) | .793 | .786 | .801 | .810 | .741 | .675 | .646 | .706 |
| GeBERTa Dada et al. (2023) | .823 | .817 | .829 | .839 | .769 | .684 | .702 | .667 | |
| GeBERTa Dada et al. (2023) | .837 | .848 | .826 | .834 | .757 | .669 | .700 | .640 | |
@@
| GBERT-BioM-Translation | .825 | .851 | .801 | .808 | .716 | .661 | .642 | .681 | |
| medical | GBERT-BioM-Translation | .833 | .860 | .807 | .811 | .714 | .692 | .677 | .707 |
@@
| GBERT-Clinical | .833 | .853 | .815 | .807 | .726 | .696 | .670 | .725 | |
| GBERT-Clinical | .843 | .876 | .812 | .806 | .710 | .716 | .692 | .742 | |
| GeBERTa-Clinical | .804 | .816 | .792 | .845 | .762 | .680 | .699 | .662 | |
| clinical | GeBERTa-Clinical | .833 | .874 | .796 | .846 | .768 | .703 | .677 | .730 |
@@
| XLM-R | .841 .003 | .823 .007 | .860 .007 | .775 | .764 | .786 | |
| GELECTRA | .850 .006 | .835 .005 | .865 .007 | .780 | .769 | .792 | |
| general | mDeBERTa V3 | .843 .005 | .824 .007 | .862 .007 | .768 | .753 | .783 |
| GeBERTa | .848 .007 | .830 .010 | .866 .007 | .772 | .761 | .783 | |
| GeBERTa | .847 .008 | .825 .003 | .872 .001 | .772 | .758 | .786 | |
@@
| GBERT-BioM-Translation | .842 .006 | .824 .007 | .861 .007 | .780 | .766 | .794 | |
| medical | GBERT-BioM-Translation | .844 .007 | .825 .007 | .864 .009 | .786 | .779 | .793 |
@@
| GBERT-Clinical | .847 .009 | .828 .009 | .867 .011 | .772 | .763 | .781 | |
| GBERT-Clinical | .844 .007 | .825 .007 | .864 .009 | .785 | .778 | .792 | |
| GeBERTa-Clinical | .846 .006 | .823 .007 | .870 .009 | .768 | .757 | .780 | |
| clinical | GeBERTa-Clinical | .855 .007 | .832 .006 | .878 .011 | .773 | .760 | .787 |
@@ 9. Discussion and Limitations
Overall, the results show that especially the addition of medical pre-training data gives a performance advantage, but not necessarily the quality of the data. For example, a 6.9 percentage points improvement in F1-Score performance was measured on the GraSCCo dataset between GBERT and GBERT-Clinical. However, the difference between GBERT-Clinical and GBERT-BioM-Translation is only 2.4 percentage points. Similarly, in RadQA the difference between GBERT and GBERT-BioM-Translation is 1.4 percentage points, but the difference between the two further pre-trained models is only 0.1 percentage points. We conclude that the presence of medical data is crucial for pre-training, but not necessarily its quality or proximity for the downstream task. The small differences between our GeBERTa-Clinical and standard GeBERTa models further support this hypothesis, as the standard version of GeBERTA already contained translations. In this context, it is also worth noting that the difference between GBERT-Clinical and GBERT-BioM-Translation models is particularly evident on the GraSCCo task, even though it consists of synthetically generated documents and not real clinical texts. In contrast, smaller differences are evident in clinical tasks such as RadQA.
@@ 13. Bibliographical References
[…]

### v2

@@ Abstract
| Ahmad Idrissi-Yaghir1,2∗, Amin Dada3∗, Henning Schäfer1,4∗, |
| Kamyar Arzideh3, Giulia Baldini3,11, Jan Trienes3, Max Hasin3, |
| Jeanette Bewersdorff5, Cynthia S. Schmidt3,4, Marie Bauer3, |
| Kaleb E. Smith6, Jiang Bian7,8, Yonghui Wu7,8, Jörg Schlötterer9,10, |
| Torsten Zesch5, Peter A. Horn4, Christin Seifert9, |
| Felix Nensa3,11, Jens Kleesiek3,12,13,14,Christoph M. Friedrich1,2 |
| 1 Department of Computer Science, University of Applied Sciences and Arts Dortmund, Dortmund, Germany |
| 2 Institute for Medical Informatics, Biometry and Epidemiology, University Hospital Essen, Essen, Germany |
| 3 Institute for AI in Medicine (IKIM), University Hospital Essen (AöR), Essen, Germany |
| 4Institute for Transfusion Medicine, University Medicine Essen, Essen, Germany |
| 5Computational Linguistics, CATALPA FernUniversität in Hagen, Germany |
| 6NVIDIA, Santa Clara, CA, USA |
| 7Department of Health Outcomes and Biomedical Informatics, College of Medicine, |
@@
| 8Cancer Informatics and eHealth core, University of Florida Health Cancer Center, |
@@
| 9University of Marburg, Marburg, Germany |
| 10University of Mannheim, Mannheim, Germany |
| 11University Hospital Essen, Institute of Interventional and Diagnostic Radiology and Neuroradiology, |
@@
| 12Cancer Research Center Cologne Essen (CCCE), West German Cancer Center Essen, |
@@
| 13German Cancer Consortium (DKTK, Partner site Essen), Heidelberg, Germany |
| 14Department of Physics, TU Dortmund, Dortmund, Germany |
@@ 1. Introduction
000∗ These authors contributed equally to this workIn recent years, pre-trained language models (PLMs) such as BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019) have become crucial in the field of natural language processing (NLP). These models have significantly enhanced the performance of a wide range of tasks, including document and token classification, as well as machine translation and text summarization. The success of these models is primarily attributed to their transformer-based architecture Vaswani et al. (2017) and their training on large amounts of unlabeled data. However, since these models are often trained on general data sources such as Wikipedia, news articles, and books, their effectiveness may be limited in specific domains, such as medicine or finance, which have distinct terminologies and writing styles. To achieve better results in these specialized fields, it is essential to use language models that are tailored to these specific domains. Building on this concept, language models can be adapted to specialized domains through two methods. The first approach involves training models from scratch on unlabeled data from the desired domain. This method ensures that the model is grounded in the unique characteristics of the target domain from the beginning. The second approach relies on continuous pre-training. Instead of starting from scratch, existing general-purpose pre-trained models can be used and refined through further pre-training on the domain-specific unlabeled data Gururangan et al. (2020). This allows a transition that shifts the focal point of the model from a broad scope to one specific to the particularities of the target domain.
@@ 4. Base Models
In this section, the models included in our benchmark are described. As a baseline for our evaluations, we use models that have been pre-trained on extensive datasets and have demonstrated strong performance on a variety of general and medical NLP tasks. The general German language models GBERTbase, GBERTlarge and GELECTRAlarge Chan et al. (2020) were primarily considered. They were trained on four different datasets: the German portion of the Open Super-large Crawled ALMAnaCH coRpus (OSCAR) (Ortiz Suárez et al., 2020), German Wikipedia dump, The Open Parallel Corpus (OPUS) (Tiedemann, 2012), and Open Legal Data (Ostendorff et al., 2020). Another model is GottBERT (Scheible et al., 2020), a German RoBERTabase model trained on the German part of OSCAR data. Furthermore, two multilingual models XLM-R Conneau et al. (2020) and mDeBERTa V3 He et al. (2023) were also considered. Additionally, GeBERTabase and GeBERTalarge Dada et al. (2023) were explored, with these DeBERTa v2 He et al. (2021) based models pre-trained from scratch on a combination of data sources ranging from Wikipedia to medical datasets. Finally, the German biomedical and clinical models BioGottBERT Lentzen et al. (2022) and medBERTde Bressem et al. (2024) were also examined to provide additional comparative insights.
@@ 5. Pre-training
Continuous pre-training of several publicly available general domain models was performed with the aforementioned datasets. First, we continued the pre-training of the GeBERTabase and GeBERTalarge models with the clinical dataset. The objective was to quantify the contribution of clinical data to the performance of the models. While the original model has already been trained on medical texts obtained mainly from MIMIC-III, translated PubMed abstracts, and filtered CC100 Wenzek et al. (2020), it has not seen any real-world German clinical data. Both models were trained for 200k AdamW Loshchilov and Hutter (2019) optimization steps with a batch size of 512. The learning rate was set to 3e-5 for the base model and 2e-5 for the large model. Additionally, the GBERTbase and GBERTlarge models were further pre-trained on the clinical data with the same parameters. These models were not explicitly pre-trained on medical texts before.
In order to separately quantify the influence of translation-based medical texts and clinical documents, additional pre-training experiments were conducted with the GBERTbase and GBERTlarge models. Both models underwent further pre-training on the translated dataset. The GBERT was trained for 73k steps with a learning rate of 5e-5 and a batch size of 144. On the other hand, the GBERT was trained for 75k steps using a batch size of 336. Due to hardware limitations, the initial experiments were conducted separately, reflecting the differences in batch sizes between the experiments. The AdamW optimizer was also employed during this pre-training for the optimization process.
@@ 6. Downstream Datasets & Tasks
The RadQA dataset Dada et al. (2024) is an extractive question-answering dataset created from 1,223 anonymized radiology reports of brain CT scans from a large hospital in Germany. For its development, three medical student assistants in their sixth and eighth semesters were assigned to annotate 29,273 question-answer pairs. The annotators were provided with a list of questions designed with the input of a radiologist. This decision was influenced by the unique nature of radiology queries and the challenges of annotating sensitive clinical data. The annotators generated one custom question for every third report to ensure variety.
@@ 8. Results
Across all tasks, our clinically pre-trained models achieved the highest F1-Scores, with the exception of GGPONC 2.0, where one of the translation models achieved the highest score. This is especially evident in the results of the GraSCCo dataset, where clinical pre-training improved the performance of GBERTbase by 5.4 percentage points in F1-Score and the performance of GBERTlarge by 6.9 percentage points. In addition, the GBERT-BioM-translation models were able to outperform the general models. For instance, pre-training on translated texts resulted in an improvement of 2.1 percentage points over GBERTbase.
Although BRONCO150 is the only NER task with real-world clinical documents, there is no observable performance difference between MedBERTde, GeBERTabase, and our clinical base models. However, the GeBERTa-Clinicallarge model outperformed the second-best model by 0.5 percentage points. The minor differences between the various models and the lack of clear advantages of clinical models can be attributed to the small size of the dataset. Interestingly, on GGPONC 2.0, various general domain models are on par with clinical models or even better. In the case of GeBERTabase, the additional clinical pre-training decreased its performance. It is worth noting that the dataset consists of guidelines for oncologists written in a different writing style than clinical documents. Overall, the translation-based models achieved the highest results.
In the multi-label classification CLEF eHealth 2019 task, large models generally performed better. For example, GBERTlarge has an F1-Score that is around 1.6 percentage points better than its base variant. Domain-specific continuous pre-training on the translations reached strong performance across all metrics. Prioritizing a balance between precision and recall favors the GELECTRAlarge model, while GeBERTa-Clinicallarge surpasses in precision. The general domain GBERTlarge model provides the highest recall, although its precision is lower compared to others.
Looking at the difference between GBERT-Clinicalbase and GBERT-BioM-Translationbase in Table 4, the latter achieves comparable results to its clinical counterpart on BRONCO, GGPONC 2.0, and RadQA. Only on GraSCCo the results of the GBERT models trained on the translations are worse. Compared to the baseline GBERT general models, all further pre-trained models have an advantage. In this case, we see no clear advantage of training on German clinical data over translated texts.
@@
| GBERTbase Chan et al. (2020) | .816 | .818 | .815 | .794 | .707 | .642 | .617 | .676 | |
| GBERTlarge Chan et al. (2020) | .832 | .802 | .865 | .809 | .718 | .647 | .617 | .680 | |
@@
| XLM-Rlarge Conneau et al. (2020) | .804 | .781 | .829 | .813 | .731 | .674 | .655 | .694 | |
| GELECTRAlarge Chan et al. (2020) | .827 | .826 | .828 | .812 | .725 | .681 | .702 | .661 | |
| general | mDeBERTa V3base He et al. (2023) | .793 | .786 | .801 | .810 | .741 | .675 | .646 | .706 |
| GeBERTabase Dada et al. (2023) | .823 | .817 | .829 | .839 | .769 | .684 | .702 | .667 | |
| GeBERTalarge Dada et al. (2023) | .837 | .848 | .826 | .834 | .757 | .669 | .700 | .640 | |
@@
| GBERT-BioM-Translationbase† | .825 | .851 | .801 | .808 | .716 | .661 | .642 | .681 | |
| medical | GBERT-BioM-Translationlarge† | .833 | .860 | .807 | .811 | .714 | .692 | .677 | .707 |
@@
| GBERT-Clinicalbase† | .833 | .853 | .815 | .807 | .726 | .696 | .670 | .725 | |
| GBERT-Clinicallarge† | .843 | .876 | .812 | .806 | .710 | .716 | .692 | .742 | |
| GeBERTa-Clinicalbase† | .804 | .816 | .792 | .845 | .762 | .680 | .699 | .662 | |
| clinical | GeBERTa-Clinicallarge† | .833 | .874 | .796 | .846 | .768 | .703 | .677 | .730 |
@@
| XLM-Rlarge | .841 .003 | .823 .007 | .860 .007 | .775 | .764 | .786 | |
| GELECTRAlarge | .850 .006 | .835 .005 | .865 .007 | .780 | .769 | .792 | |
| general | mDeBERTa V3base | .843 .005 | .824 .007 | .862 .007 | .768 | .753 | .783 |
| GeBERTabase | .848 .007 | .830 .010 | .866 .007 | .772 | .761 | .783 | |
| GeBERTalarge | .847 .008 | .825 .003 | .872 .001 | .772 | .758 | .786 | |
@@
| GBERT-BioM-Translationbase† | .842 .006 | .824 .007 | .861 .007 | .780 | .766 | .794 | |
| medical | GBERT-BioM-Translationlarge† | .844 .007 | .825 .007 | .864 .009 | .786 | .779 | .793 |
@@
| GBERT-Clinicalbase† | .847 .009 | .828 .009 | .867 .011 | .772 | .763 | .781 | |
| GBERT-Clinicallarge† | .844 .007 | .825 .007 | .864 .009 | .785 | .778 | .792 | |
| GeBERTa-Clinicalbase† | .846 .006 | .823 .007 | .870 .009 | .768 | .757 | .780 | |
| clinical | GeBERTa-Clinicallarge† | .855 .007 | .832 .006 | .878 .011 | .773 | .760 | .787 |
@@ 9. Discussion and Limitations
Overall, the results show that especially the addition of medical pre-training data gives a performance advantage, but not necessarily the quality of the data. For example, a 6.9 percentage points improvement in F1-Score performance was measured on the GraSCCo dataset between GBERTlarge and GBERT-Clinicallarge. However, the difference between GBERT-Clinicallarge and GBERT-BioM-Translationlarge is only 2.4 percentage points. Similarly, in RadQA the difference between GBERTbase and GBERT-BioM-Translationbase is 1.4 percentage points, but the difference between the two further pre-trained models is only 0.1 percentage points. We conclude that the presence of medical data is crucial for pre-training, but not necessarily its quality or proximity for the downstream task. The small differences between our GeBERTa-Clinical and standard GeBERTa models further support this hypothesis, as the standard version of GeBERTA already contained translations. In this context, it is also worth noting that the difference between GBERT-Clinical and GBERT-BioM-Translation models is particularly evident on the GraSCCo task, even though it consists of synthetically generated documents and not real clinical texts. In contrast, smaller differences are evident in clinical tasks such as RadQA.
@@ 13. Bibliographical References
- Alsentzer et al. (2019) Emily Alsentzer, John Murphy, William Boag, Wei-Hung Weng, Di Jindi, Tristan Naumann, and Matthew McDermott. 2019. Publicly available clinical BERT embeddings. In Proceedings of the 2nd Clinical Natural Language Processing Workshop, pages 72–78, Minneapolis, Minnesota, USA. Association for Computational Linguistics.
@@
- Dada et al. (2024) Amin Dada, Tim Leon Ufer, Moon Kim, Max Hasin, Nicola Spieker, Michael Forsting, Felix Nensa, Jan Egger, and Jens Kleesiek. 2024. Information extraction from weakly structured radiological reports with natural language queries. European Radiology, 34(1):330–337.